1. Ten years ago, the Circus Corp. sold a 20-year bond issue with a 9 percent annual coupon rate and a 3 percent call premium. Today, Circus called the bonds. The bonds originally were sold at their face value of $1,000. Compute the realized rate of return (yield to call) for investors who purchased the bonds when they were issued and who surrender them today in exchange for the call price.

2.   (EXCEL TEMPLATE) Grass Whacker’s is considering whether or not to refund a $90 million, 6 percent coupon, 30 year bond issue that was sold 8 years ago. It is amortizing $1.8 million of flotation costs over the issue’s 30-year life. A new 22-year issue would carry an interest rate of 5.35 percent. A call premium of 5 percent would be required to retire the old bonds and flotation costs on the new issue would be $1.35 million (also to be amortized). Grass Whacker’s marginal tax rate is 34 percent. The new bonds would be issued 1 month before the old bonds are called, with the proceeds being invested in short-term securities returning 1.5 percent annually. What is the NPV of the refund? Use the NPV function in Excel and the PV function in Excel. Answers should be the same.

Use the Excel Template and the following information to answer the next three questions:

Peterson Packaging Inc. does not currently pay dividends. The company will start with a $0.50 dividend at the end of year three and grow it by 10% for each of the next seven years until it nearly reaches $1.00. After seven years of growth, it will fix its dividend at $1.00 forever. The required return on the stock is 15%.

3. (EXCEL TEMPLATE) What is the current stock price? Use the NPV function.

4. (EXCEL TEMPLATE) What is the expected stock price in years 1-10? Use the NPV function.

5. (EXCEL TEMPLATE) Calculate the dividend yield and capital gains yield that an investor should expect for each year.

Use the financial statements shown below to answer the next three questions. Free cash flow is expected to grow at 4 percent after 2019. The weighted average cost of capital is 6.9 percent. The bonds are currently selling at 103% of par. The preferred stock has a current market value of $51.51 million.

Balance Sheet (in millions)

Actual

2018

Projected

2019

Actual

2018

Projected

2019

Cash

81.90

61.60

Accounts payable

26.10

22.20

Marketable securities

53.80

54.20

Notes payable

101.00

4.70

Accounts Receivable

26.40

27.00

Accruals

8.00

7.37

Inventory

198.20

186.90

Total current liabilities

135.10

34.27

Total Current Assets

360.30

329.70

Long term bonds

9.20

4.50

Preferred Stock

50.20

52.40

Common stock

230.00

230.00

Retained earnings

307.00

367.23

Net fixed assets

371.20

358.70

Total common equity

537.00

597.23

Total assets

731.50

688.40

Total liabilities & equity

731.50

688.40

Income Statement (in millions)

Actual 2018

Projected 2019

Sales

907.30

955.10

Operating expenses

764.40

802.80

Depreciation

21.30

21.30

Earnings before interest & taxes

121.90

131.00

Interest

0.60

0.60

Earnings before taxes

121.30

130.40

Taxes

45.90

50.10

Net income before preferred dividends

75.40

80.30

Preferred dividends

2.50

1.60

Net Income available for common

72.90

78.70

Common dividends

16.40

18.47

Addition to retained earnings

56.50

60.23

Number of shares (in millions)

12.70

12.70

6. What is the free cash flow for 2019?

7. What is the value of operations as of 2018?

8. What is the price per share for 2018?
Answer #1

1)

A1 B C D E F G H I J K L M N O
2
3 Face Value $1,000
4 Issue Price $1,000
5 Coupon Rate 9%
6 Call Premium 3%
7 Number of years before bond was called 10 Years
8 Annual Coupon $90 =D3*D5
9 Since call premium is 3%, amount paid by company at the time of calling the bond is $1,000*(1+3%) = $1,030 for each bond.
10
11 Cash Flow to investor is
12 Year 0 1 2 3 4 5 6 7 8 9 10
13 Price paid to purchase the bond ($1,000)
14 Coupon Received $90 $90 $90 $90 $90 $90 $90 $90 $90 $90
15 Amount received when bond was recalled $1,030
16 Total Cash Flow to investor ($1,000) $90 $90 $90 $90 $90 $90 $90 $90 $90 $1,120 =N13+N14+N15
17
18 Realized return to investor will be the rate at which Net Present Value of above cash flow will be zero.
19 This can be calculated using IRR function as follows:
20 Realized rate of return 9.20% =IRR(D16:N16)
21
22 Hence realized rate of return to the investor is 9.20%
23

Formula sheet

A1 B C D E F G H I J K L M N O
2
3 Face Value 1000
4 Issue Price 1000
5 Coupon Rate 0.09
6 Call Premium 0.03
7 Number of years before bond was called 10 Years
8 Annual Coupon =D3*D5 =D3*D5
9 Since call premium is 3%, amount paid by company at the time of calling the bond is $1,000*(1+3%) = $1,030 for each bond.
10
11 Cash Flow to investor is
12 Year 0 =D12+1 =E12+1 =F12+1 =G12+1 =H12+1 =I12+1 =J12+1 =K12+1 =L12+1 =M12+1
13 Price paid to purchase the bond =-D4
14 Coupon Received =$D$8 =$D$8 =$D$8 =$D$8 =$D$8 =$D$8 =$D$8 =$D$8 =$D$8 =$D$8
15 Amount received when bond was recalled 1030
16 Total Cash Flow to investor =D13+D14+D15 =E13+E14+E15 =F13+F14+F15 =G13+G14+G15 =H13+H14+H15 =I13+I14+I15 =J13+J14+J15 =K13+K14+K15 =L13+L14+L15 =M13+M14+M15 =N13+N14+N15 =N13+N14+N15
17
18 Realized return to investor will be the rate at which Net Present Value of above cash flow will be zero.
19 This can be calculated using IRR function as follows:
20 Realized rate of return =IRR(D16:N16) =IRR(D16:N16)
21
22 Hence realized rate of return to the investor is =D20
23